# DriveSOTIF: Advancing Perception SOTIF Through Multimodal Large Language Models

### v1

@@ Abstract
Human drivers naturally possess the ability to perceive driving scenarios, predict potential hazards, and react instinctively due to their spatial and causal intelligence, which allows them to perceive, understand, predict, and interact with the 3D world both spatially and temporally. Autonomous vehicles, however, lack these capabilities, leading to challenges in effectively managing perception-related Safety of the Intended Functionality (SOTIF) risks, particularly in complex and unpredictable driving conditions. To address this gap, we propose an approach that fine-tunes multimodal language models (MLLMs) on a customized dataset specifically designed to capture perception-related SOTIF scenarios. Model benchmarking demonstrates that this tailored dataset enables the models to better understand and respond to these complex driving situations. Additionally, in real-world case studies, the proposed method correctly handles challenging scenarios that even human drivers may find difficult. Real-time performance tests further indicate the potential for the models to operate efficiently in live driving environments. This approach, along with the dataset generation pipeline, shows significant promise for improving the identification, cognition, prediction, and reaction to SOTIF-related risks in autonomous driving systems. The dataset and information are available: https://github.com/s95huang/DriveSOTIF.git
@@ I Introduction
In autonomous driving (AD), safety can be classified into functional safety and Safety of the Intended Functionality (SOTIF). Functional safety addresses issues with hardware and software related to a specific function, ensuring that the system operates correctly in response to its inputs. SOTIF, on the other hand, addresses safety concerns that arise not from system failures but from situations where the system performs as intended but may still lead to unsafe conditions due to external factors or system limitations [1].
@@
Perception systems in autonomous vehicles (AVs), which are tasked with detecting, classifying, and predicting based on environmental stimuli, are particularly vulnerable to SOTIF-related challenges. Issues such as sensor limitations in adverse weather, difficulty in recognizing unusual road users, or interpreting rare and complex scenarios can lead to unsafe decisions [2]. The dynamic and unpredictable nature of real-world driving adds complexity, as sudden obstacles or erratic behavior from other road users can easily overwhelm current algorithms [3].
@@
Detecting and addressing SOTIF risks are therefore essential for the safe operation of AVs. Existing methods adopt rule-based methods to quantify perception-related risks by measuring entropy levels and classifying them into specific categories [3, 4]. In contrast, human drivers can naturally assess the risk and uncertainty with certain objects when driving. This capability of assessing SOTIF-related risk and selecting the best action based on personal experience, knowledge, common sense, etc., remains absent from existing literature.
Recent advances in multimodal large language models (MLLMs) offer new opportunities to improve the understanding and management of SOTIF risks. By combining visual perception with language-based reasoning, these models can mimic the human ability to assess risk and uncertainty in driving scenarios.
@@ II-A Perception-related SOTIF
Perception SOTIF risks arise when an autonomous vehicle’s perception system functions correctly at the hardware and software levels but still fails to accurately interpret the driving environment due to external factors [3]. Since perception is fundamental for downstream decision-making and planning, any misinterpretation can propagate through the system, leading to unintended emergency braking, unsafe lane changes, or even collisions. These risks are further amplified by factors like sensor degradation due to rain, fog, snow, and glare, and uncommon object appearances [4]. Identifying and mitigating perception-related SOTIF risks is crucial for maintaining vehicle safety, particularly when an autonomous vehicle operates outside its predefined Operational Design Domain (ODD) [5, 6].
Current approaches to detect and mitigate perception-related SOTIF risks can be classified into rule-based and learning-based methods. Rule-based approaches, such as Fault Tree Analysis (FTA) [7], System Theoretic Process Analysis (STPA) [8, 9], and formal verification [10], attempt to identify perception-triggering conditions and system failures through structured models [9]. However, these methods struggle with dynamic and unpredictable environments, limiting their applicability to real-world driving conditions [11].
In contrast, learning-based and probabilistic uncertainty quantification techniques aim to detect and mitigate perception-related SOTIF risks through data-driven models. Methods such as Bayesian Risk Decomposition [12], deep ensemble [4, 3, 13], and sensor fusion techniques have been explored to quantify uncertainty in perception outputs and improve reliability [7]. Despite these advancements, learning-based methods often suffer from black-box behavior, making it challenging to interpret the underlying process, and poor generalization to edge cases, particularly in long-tail traffic scenarios.
II-B LLM in AD
Early research in the application of LLMs for autonomous driving primarily focused on enhancing interaction and decision-making (DM) through text-based LLMs. For instance, Cui et al. introduced the DriveLLM framework and demonstrated the feasibility of LLM-enhanced decision-making through experimental studies [14]. By integrating visual feature encoders with text-based LLMs, these models gain the ability to interpret visual inputs [15, 16, 17]. This combination allows the model to process and understand both visual and linguistic data, enabling more sophisticated reasoning and decision-making capabilities [18, 19].
In AD, this integration enhances the system’s capacity to interpret complex environments and respond to visual stimuli by combining visual inputs with language-based reasoning. Additionally, by combining image and textual information from other modules, such as vehicle states, safety, decision-making, planning, and control, the system can make more informed responses and better understand and adapt to dynamic driving scenarios [14, 20, 21]. This approach makes MLLMs well-suited for handling real-world complex driving scenarios, allowing autonomous systems to adjust to unpredictable conditions and improve safety and reliability [22, 23, 24, 25, 19, 20].
However, much of the existing research in the AD field relies on proprietary models, which require continuous access to internet. For open-source models, they are usually trained on general-purpose datasets like COCO Caption [26], VQAv2 [27], making them less suitable for AD domain applications.
One major issue with existing LLM-related autonomous driving datasets is the lack of diversity in driving scenarios. Datasets like DRAMA[28] and nuScenes-QA[23] are comprehensive in size but tend to focus on specific environments. Models trained on these datasets may face difficulties when geographical domain transfer or road conditions happen. Another significant limitation is the complexity and cost of annotations. Datasets like NuScenes-QA [23], Refer-KITTI [29] adopt a markup-based approach to generate template-based questions and answers. While this method balances well between dataset quality and cost, it results in a dataset that is not human-like. This limits the model’s ability to handle more open-ended or complex tasks that require a broader range of responses. Lastly, while datasets like DRAMA [28] emphasize safety-critical events, they don’t fully capture the complexity of real-world hazardous driving situations. In actual driving, safety risks come from unpredictable factors like sudden obstacles, changing weather, or erratic behavior from other road users. Current datasets often miss these scenarios, making it harder for models to perceive, understand, and respond to risks in real-time. Addressing this gap is essential for ensuring the safety and reliability of Level 3-5 autonomous vehicles.
III Overview
III-A Motivation and Contributions
Based on the literature review, despite the advancements in rule-based and learning-based approaches for detecting perception-related SOTIF risks, existing methods struggle with adaptability, interpretability, and robustness in dynamic driving environments. Rule-based techniques rely on pre-defined conditions that fail to generalize to unpredictable real-world scenarios, while deep learning-based methods often suffer from black-box behavior, poor generalization to edge cases, and vulnerability to adversarial conditions. Most importantly, these models lack contextual reasoning capabilities, making it difficult to understand, predict, and react appropriately to safety-critical situations. These limitations highlight the need for a novel approach that can detect and understand perception-related SOTIF risks.
Drawing inspiration from the way human drivers naturally perceive, reason, and act in complex environments, we propose leveraging MLLMs to address perception-related SOTIF risks. We fine-tune pre-trained MLLMs on DriveSOTIF, a domain-specific dataset carefully designed to capture the nuances of safety-critical driving scenarios. Through supervised fine-tuning (SFT), we enhance pre-trained MLLMs with domain-specific safety knowledge, allowing them to detect, reason, and respond effectively to safety-critical driving situations. This approach aims to enhance situational awareness, risk assessment, and real-time decision-making, leading to safer and more reliable autonomous vehicle systems. The key contributions of this study are summarized as follows:
  1. 1.
    Integration of MLLMs into SOTIF Research: We evaluate pre-trained MLLMs and enhance their performance through SFT on a custom domain-specific dataset. This comprehensive process—from dataset creation to model deployment—significantly improves the model’s ability to detect, reason about, and mitigate perception-related SOTIF risks in autonomous driving.
  2. 2.
    DriveSOTIF Dataset: The DriveSOTIF dataset is introduced as the first VQA and image captioning dataset specifically tailored for perception-related SOTIF risks. Its dataset generation pipeline can be applied to other autonomous driving tasks, making MLLM dataset creation more cost-effective and less prone to bias.
  3. 3.
III-B Methodology
Figure 1: Supervised fine-tuneing of pre-trained multimodal LLMs on domain-specific dataset for enhanced perception SOTIF cognition and reaction
As illustrated in Figure 1, the proposed approach consists of three main stages:
  1. 1.
    Domain-Specific Data Curation (DriveSOTIF): We first constructed the DriveSOTIF datasetto to captures diverse, safety-critical driving scenarios, ranging from adverse weather conditions to rare object encounters. Each scenario is represented through images, visual-question answering (VQA) prompts, and image-captioning annotations that focus on perception-related SOTIF risks. The DriveSOTIF dataset is generated using a pipeline of collaborative MLLM agents, designed to generate the desired dataset efficievely with reduced hallucinatnions and bias. A round of human review is performed to ensure dataset quality and overall reliability.
  2. 2.
    Supervised Fine-Tuning: We begin with pre-trained MLLMs and fine-tune them on the DriveSOTIF dataset through a SFT procedure. We employ Low-Rank Adaptation (LoRA) and quantized LoRA (QLoRA), which allow us to a small portion of the model rather than retraining the entire model. This strategy significantly reduces computational overhead and memory usage, which is more practical for real-world applications.
  3. 3.
    Deployment and Validation: After fine-tuning, we benchmark the fine-tuned models on the test split of the DriveSOTIF dataset and compare their performance against the original pre-trained MLLMs. We then select two lightweight, high-performing models to examine how different model sizes and quantization levels affect real-time inference. Finally, we validate these models in real-world settings, including university campuses, city streets, and proprietary dashcam footage collected in China.
IV DriveSOTIF Dataset
IV-A Image Selection
@@
Today’s passenger vehicles often contain multiple cameras, posing a challenge for VQA tools that need to incorporate these inputs to effectively assist in BEV-based perception SOTIF problems. For a dataset newly introduced to SOTIF, we aim to start with simple one-image and one-question pairs. This approach allows for comparisons with existing VQA datasets and algorithms and encourages further research in this area.
TABLE I: DriveSOTIF Dataset Answer and Question Types
@@
Open-ended No fixed question structure and have multiple correct answers
@@
Uncertainty Uncertainty level with a certain object (Low, Medium, High)
Existence Does an object of class XXX exist?
Type What is type of object at the bottom left?
Count How many XXX are there?
Key object Is this object a key object?
For the VQA subtask, we designed our dataset such that it supports close-ended and open-ended Q&A, as shown in Table I. Closed-ended questions are divided into five categories: uncertainty, existence, type, counting, and key object. These questions typically use templates or examples for generation, providing straightforward answers with limited choices such as yes/no or other specific options.
Open-ended Q&A are designed for higher levels of reasoning, such as causal inference and complex reasoning. These questions do not follow a fixed structure and can have multiple correct answers, offering more flexibility and depth in responses. These questions are focused on SOTIF risk identification, explanation, and providing recommended actions.
To ensure a balanced representation of question types, we extend our dataset to include ten questions per image: five closed-ended and five open-ended.
This approach ensures a diverse set of questions and promotes a comprehensive evaluation of the dataset.
IV-B Dataset Generation through Collaborative LLM-Agents.
To automatically generate the VQA and image captioning datasets, we leverage a multi-agent framework utilizing proprietary LLMs with vision capabilities. In this study, we use OpenAI’s GPT-4v, GPT-4o, and Claude 3 Opus from Anthropic. The dataset generation pipeline, as illustrated in Figure 2, is divided into distinct roles for image captioning, question generation, answer generation, and validation/post-processing.
@@
Image Captioning Agent: The image captioning agent aims to generate the caption for the image based on a prompt. Importantly, we alternates between two LLM backends for question generation and captioning. This approach is adopted to minimize bias that might arise from relying on a single proprietary LLM model. Sample images and captions are provided in Table II.
Question Generation Agent: The primary objective of this agent is to generate natural, human-like questions based on the input query, captions generated, along with 2D object detection and uncertainty information  [^1] Please refer to supplemental material for object detection and uncertainty quantification. . Specifically, the question-generating agent is instructed to generate questions with varying difficulties and types. For closed-end questions, the agent randomly selects the question type from Table I. For open-ended questions, the agent focuses on advanced QA like causal inference, including identifying and explaining the SOTIF risk and providing recommended action to prevent the risk, if possible. Sample questions generated are shown in Table II.
TABLE II: Sample labeled Questions & Answers and image captions from SOTIF-VQA dataset.
@@
Answer Generation Agent: In the VQA dataset format, each question is answered by ten responses from ten annotators. While this approach ensures accuracy, it is both costly and time-consuming. To streamline this process, the answer generation agent is designed to provide ten answers for each question, given an image. To achieve this, we employ a model aggregation method that utilizes GPT-4V and Claude3-Opus. For each image, the image file and associated questions are fed into the LLMs. Instead of sequentially querying a single model multiple times, the system generate five answers for each question in a single pass through each model. This method is more efficient and cost-effective compared to sending the questions individually.
@@
Validation and Post-processing Agent: As shown in Figure 2, newly generated answers are then sent to the validation and post-processing agent, which contains three post-processing mechanisms. The first and second mechanisms are to ensure that captions and questions generated are related to the image and perception SOTIF context. The third mechanism is to ensure that answers generated are correct based on the input image and question. If any inconsistencies are found, the agent regenerates the captions, questions or answers as needed. This validation process leverages OpenAI’s GPT4o as the LLM backend to ensure high-quality outputs that align with the perception SOTIF context.
IV-C Dataset Overview
The multi-agent system is implemented to generate the dataset using different temperature settings to ensure a variety of questions/answers and captions. After the dataset generation, the data entries are inspected to ensure correct structure and then organized and converted into the desired format. To ensure dataset quality and effectiveness, we conduct a round of human review and verification. This manual process identifies and corrects errors that automated systems might miss or generate.
Our dataset contains 1114 images, 1114 captions, 5570 questions, and 55700 answers  [^2] It should be noticed that collecting SOTIF-related scenarios is partially challenging due to their rarity and unpredictability in real-world settings, resulting in a relatively small dataset size. However, we are actively working with PeSOTIF authors to expand the dataset by adding new scenarios.. The image captioning dataset is organized into the COCO caption [26] format, while the VQA dataset is organized using VQAV2 [27] format [^3] For compatibility with recent research, we converted the image-captioning and VQA dataset into the LLaVA format. We also generated a Chinese version of the dataset by translating the original dataset using GPT-4o.. These datasets are split into train/validation and test sets at a ratio of 0.8 and 0.2, as shown in Table III.
TABLE III: Training and Validation split statistics.

### v2

# DriveSOTIF: Advancing SOTIF Through Multimodal Large Language Models

@@ Abstract
Human drivers possess spatial and causal intelligence, enabling them to perceive driving scenarios, anticipate hazards, and react to dynamic environments. In contrast, autonomous vehicles lack these abilities, making it challenging to manage perception-related Safety of the Intended Functionality (SOTIF) risks, especially under complex or unpredictable driving conditions. To address this gap, we propose fine-tuning multimodal large language models (MLLMs) on a customized dataset specifically designed to capture perception-related SOTIF scenarios. Benchmarking results show that fine-tuned MLLMs achieve an 11.8% improvement in close-ended VQA accuracy and a 12.0% increase in open-ended VQA scores compared to baseline models, while maintaining real-time performance with a 0.59-second average inference time per image. We validate our approach through real-world case studies in Canada and China, where fine-tuned models correctly identify safety risks that challenge even experienced human drivers. This work represents the first application of domain-specific MLLM fine-tuning for SOTIF domain in autonomous driving. The dataset and related resources are available at github.com/s95huang/DriveSOTIF.git
@@ I Introduction
In autonomous driving (AD), safety is commonly classified into functional safety and Safety of the Intended Functionality (SOTIF). Functional safety concerns failures in hardware or software that result in unsafe operation. In contrast, SOTIF addresses hazards that occur not due to malfunctions, but when the system operates as intended yet produces unsafe outcomes because of external factors or inherent limitations [1].
@@
Perception systems in autonomous vehicles (AVs),which are responsible for detecting, classifying, and predicting based on environmental inputs, are especially vulnerable to SOTIF-related risks. Factors such as sensor limitations in adverse weather, difficulty recognizing unusual road users, or rare and complex scenarios can result in unsafe decisions [2]. The dynamic, unpredictable nature of real-world driving further complicates matters: sudden obstacles or erratic behavior by other road users can easily overwhelm current algorithms [3].
@@
Identifying and mitigating SOTIF risks is therefore essential for safe AV operation. Current methods primarily use rule-based approaches to quantify perception-related risks, for example by measuring output entropy and classifying risks into predefined categories [4]. Learning-based methods, on the other hand, suffer from black-box behavior, poor generalization to edge cases, and vulnerability to adversarial conditions. In contrast, human drivers intuitively assess risk and uncertainty for specific objects as they drive, drawing on personal experience, knowledge, and common sense—capabilities that remain absent from existing automated systems.
Recent advances in multimodal large language models (MLLMs) provide new tools to address these cognitive gaps. MLLMs offer three distinct advantages over traditional approaches: (1) open-world cognition that leverages pre-trained knowledge to interpret novel scenarios; (2) causal reasoning that traces how environmental factors cascade into safety risks; and (3) contextual inference that interprets social dynamics and behavioral patterns beyond conventional algorithmic reach [5, 6]. For example, an MLLM can infer that sun glare may obscure lane markings, recognize that people gathering with umbrellas likely indicates a bus stop, or understand how rain, construction, and nighttime conditions interact to create heightened safety risks. Crucially, MLLMs provide explainable risk reasoning via natural language, articulating why scenarios pose specific risks and recommending actionable responses—capabilities essential for safety validation and user trust.
TABLE I: Perception-Related SOTIF Datasets
Dataset Year Size Image LiDAR Text Description
PeSOTIF[7] 2022 1126 frames 2D object detection dataset for long-tail traffic scenarios with diverse weather conditions and uncommon object appearances.
[…]
Motivated by human drivers’ cognitive abilities, we propose a novel framework that leverages MLLMs to address perception-related SOTIF risks. Our approach fine-tunes pre-trained MLLMs on DriveSOTIF, a domain-specific dataset carefully designed to capture safety-critical driving scenarios. Through supervised fine-tuning (SFT), we enhance these models with specialized knowledge that allows them to detect, reason about, and effectively respond to safety-critical driving situations. This approach directly improves situational awareness, risk assessment, and decision-making for safer and more reliable autonomous vehicles. The key contributions of this study are summarized as follows:
  1. 1.
    Novel Integration of MLLMs into SOTIF Research: We introduce a comprehensive approach that integrates pre-trained MLLMs with supervised fine-tuning (SFT) on a tailored, domain-specific dataset. Our method enhances the capabilities of these models to accurately detect, reason about, and effectively mitigate perception-related SOTIF risks.
  2. 2.
    DriveSOTIF Dataset: We introduce DriveSOTIF, the first visual question answering and image captioning dataset for perception-related SOTIF challenges. Our novel multi-agent generation pipeline reduces annotation costs while maintaining high quality, providing a scalable approach for developing MLLM datasets for autonomous driving.
  3. 3.
@@ II-A Perception-Related SOTIF
Perception-related SOTIF risks occur when an autonomous vehicle’s perception system functions correctly at both the hardware and software levels yet fails to accurately interpret the environment due to external factors [3]. Because perception underpins downstream tasks such as decision-making and planning, errors at this stage can propagate, leading to hazardous outcomes including false emergency braking, unsafe lane changes, or collisions. These challenges are particularly acute in adverse conditions—rain, snow, fog, glare—or in the presence of rare or unusual objects [7], especially when vehicles operate outside their defined Operational Design Domain (ODD) [11, 12].
Mitigation approaches can be broadly categorized as rule-based or learning-based. Rule-based methods—such as Fault Tree Analysis (FTA) [4], System-Theoretic Process Analysis (STPA) [13, 14], and formal verification [15]—use structured frameworks to model how specific failures can lead to unsafe behaviors. While effective for high-level system modeling, these methods often fail to capture the unpredictable and dynamic aspects of real-world environments [16].
Learning-based and probabilistic techniques focus on modeling perception uncertainty using data-driven methods. Bayesian Risk Decomposition [17], deep ensembles [7, 3, 18], and sensor fusion have shown promise in quantifying and mitigating uncertainty in perception outputs [4]. However, these models typically function as black boxes, limiting interpretability and making it difficult to diagnose failure modes. Their performance also tends to degrade in rare or edge-case scenarios that fall outside typical training distributions.
II-B MLLMs in Autonomous Driving and Dataset Challenges
Multimodal large language models (MLLMs) are attracting increasing attention in autonomous driving due to their ability to jointly process visual and textual information [19]. This enables the integration of visual inputs—such as camera feeds—with map data, vehicle state, and textual cues, leading to a more holistic understanding of the driving environment [20, 21].
From a SOTIF perspective, MLLMs are especially promising for addressing perception limitations. Their cross-modal reasoning allows them to compensate when one sensor stream is degraded—for example, when camera images are obscured by glare or LiDAR signals are weakened by fog. By integrating complementary sensory inputs, MLLMs can infer scene context, improving situational awareness and reducing the risk of misinterpretation under challenging conditions [22, 21, 23].
Figure 1: Overview of the proposed methodology for supervised fine-tuning of pre-trained MLLMs on DriveSOTIF dataset.
However, the effective deployment of MLLMs in SOTIF applications is currently hampered by dataset limitations. Most available SOTIF-related datasets, summarized in Table I, are focused on object detection tasks. For example, PeSOTIF [7] and CODA [8] target 2D detection in rare and long-tail scenarios, while SOTIF-PCOD [9] provides synthetic LiDAR point clouds. Of these, only CODA-LM [10] includes textual annotations for vision-language models, but its primary focus is not direct SOTIF risk assessment.
In addition, existing MLLM-focused driving datasets present further challenges. Datasets like DRAMA [24] and NuScenes-QA [25], while large and well-annotated, are geographically or scenically constrained, which limits their ability to generalize to new environments. Many also rely on template-based question-answer generation [25, 26], resulting in interactions that lack the nuance and diversity of natural human queries. This restricts the models’ capacity for open-ended reasoning and complex situational judgment, which are essential for safety-critical autonomous driving applications.
III DriveSOTIF
We introduce DriveSOTIF, a domain-specific dataset designed to capture perception-related SOTIF scenarios for autonomous driving. As illustrated in Figure 1, our approach combines dataset generation with supervised fine-tuning of pre-trained MLLMs.
III-A Image Selection
@@
Today’s passenger vehicles often contain multiple cameras, posing a challenge for VQA tools that need to incorporate these inputs to effectively assist in BEV-based perception SOTIF problems. For a dataset newly introduced to SOTIF, we begin with simple image-question pairs. This approach allows for comparisons with existing VQA datasets and algorithms and encourages further research in this area.
TABLE II: DriveSOTIF Dataset Answer and Question Types
@@
Open-ended No fixed question structure
@@
Uncertainty Uncertainty level with a certain object (Low, Med, High)
Existence Does an object of class (class) exist?
Type What is the type of (object) at the bottom left?
Counting How many (objects) are there?
Key object Is this (object) a key object?
@@
For the VQA subtask, we designed our dataset such that it supports close-ended and open-ended Q&A, as shown in Table II. Closed-ended questions are divided into five categories: uncertainty, existence, type, counting, and key object. These questions typically use templates or examples for generation, providing straightforward answers with limited choices such as yes/no or other specific options.
Open-ended Q&A target higher-level reasoning, including causal inference and scenario explanation. These questions do not follow a fixed structure and can have multiple correct answers, offering more flexibility and depth in responses. These questions are focused on SOTIF risk identification, explanation, and providing recommended actions.
To ensure a balanced representation of question types, we include five questions per image, with 2–3 closed-ended and 2–3 open-ended. This approach ensures a diverse set of questions and promotes a comprehensive evaluation of the dataset.
III-B Dataset Generation through Collaborative LLM-Agents.
To automatically generate the VQA and image captioning datasets, we leverage a multi-agent framework using vision-capable large language models (LLMs), specifically OpenAI’s GPT-4v, GPT-4o, and Anthropic’s Claude 3 Opus. The dataset generation pipeline, illustrated in Figure 2, includes distinct roles for image captioning, question generation, answer generation, and validation/post-processing. All prompts used in the dataset generation and validation process are provided in S.III of the supplemental material.
Image Captioning Agent: The image captioning agent aims to generate captions based on input prompts. Importantly, we alternate between two LLM backends for question generation and captioning. This approach is adopted to minimize bias that might arise from relying on a single proprietary LLM model. Sample images and captions are provided in Table III.
TABLE III: Sample labeled Questions & Answers and image captions from SOTIF-VQA dataset.
@@
Question Generation Agent: The primary objective of this agent is to generate natural, human-like questions based on the input query and captions generated, along with 2D object detection and uncertainty information  [^1] Please refer to supplemental material for object detection and uncertaintyquantification. . Specifically, the question-generating agent is instructed to generate questions with varying difficulties and types. For closed-end questions, the agent randomly selects the question type from Table II. For open-ended questions, the agent focuses on advanced QA like causal inference, including identifying and explaining the SOTIF risk and providing recommended action to prevent the risk, if possible. At the same time, the agent also generates meta data about the question, such as the question type, difficulty level, and expected answer type. Sample questions generated are shown in Table III.
@@
Answer Generation Agent: In the VQA dataset format, each question is answered by ten responses from ten annotators. While this approach ensures accuracy, it is both costly and time-consuming. To streamline this process, the answer generation agent is designed to provide ten answers for each question, given an image. To achieve this, we employ a model aggregation method that utilizes GPT-4o and Claude3-Opus. For each image, the image file and associated questions are fed into the LLMs. Instead of sequentially querying a single model multiple times, the system generates five answers for each question in a single pass through each model. This method is more efficient and cost-effective compared to sending the questions individually. Similarly, the answer generation agent outputs metadata for each answer, including the confidence score and answer type.
@@
Validation and Post-processing Agent: As shown in Figure 2, newly generated answers are then sent to the validation and post-processing agent, which contains three post-processing mechanisms. The first and second mechanisms are to ensure that captions and questions generated are related to the image and perception SOTIF context. The third mechanism is to ensure that answers generated are correct based on the input image and question. If any inconsistencies are found, the agent regenerates the captions, questions or answers as needed. This validation process leverages OpenAI’s GPT4.1 as the LLM backend to ensure high-quality outputs that align with the perception SOTIF context.
III-C Dataset Overview
The multi-agent system employed varying temperature settings during generation to ensure diversity across questions, answers, and captions. Following automated generation, all data entries underwent structural validation and format standardization, with close-ended and open-ended questions automatically categorized using metadata from the generation process. A human review of 595 samples, selected using finite population sampling [^2] More details on the finite population sampling method can be found in S.IV of the supplemental material.., indicated a 3-4% error rate, which is acceptable for automated pipelines. An additional round of manual review addressed edge cases and subtle inconsistencies that automated systems might miss or generate.
The final DriveSOTIF dataset contains 1,114 images, 1,114 captions, and 5,570 question-answer pairs  [^3] Collecting SOTIF-related scenarios is challenging due to their rarity and unpredictability in real-world settings. We are actively collaborating with PeSOTIF authors to expand the dataset with new scenarios and modalities.. The dataset adopts COCO Caption format [27] for image descriptions and VQAv2 format [28] for question-answering tasks. We partition the data using a 90/10 train-test split, as detailed in Table IV.
Table III demonstrates the dataset’s strong alignment with perception-related SOTIF challenges. Quantitatively, caption length averages 18-19 words compared to 10.5 words in COCO Caption, providing substantially richer contextual information essential for understanding complex driving scenarios. Question difficulty maintains excellent balance across easy (39.9%), medium (40.2%), and hard (19.9%) categories. Similarly, question formats split evenly between close-ended (45.7%) and open-ended (54.3%) types, which facilitates assessment of both factual accuracy and reasoning capabilities. Answer types span analysis (36.1%), yes/no and multiple-choice (28.1%), recommendations (17.8%), counting (17.0%), and identification (1.0%) tasks. This diversity enables comprehensive benchmarking of models’ ability to handle both fundamental perception tasks and higher-level causal reasoning and safety recommendations specific to SOTIF.
The generation statistics demonstrate the effectiveness of the multi-agent approach and validate our quality control mechanisms. Caption generation achieved 90.1% first-attempt success, reflecting the straightforward nature of descriptive tasks. Question generation followed with 82.7% success, while answer generation required the most refinement, averaging 1.72 attempts per item due to the complexity of ensuring factual accuracy in analytical responses. Despite this iterative process, 71.0% of answers passed initial validation, with 21.4% requiring single revisions and only 7.6% needing multiple corrections. These results demonstrate that the pipeline consistently produces high-quality, diverse, and contextually appropriate annotations for perception-related SOTIF scenarios.
TABLE IV: Dataset statistics and generation information
Dataset Split Statistics